### v2

# Benchmarking 3D Human Pose Estimation Models under Occlusions

Abstract
Human Pose Estimation (HPE) involves detecting and localizing keypoints on the human body from visual data. In 3D HPE, occlusions, where parts of the body are not visible in the image, pose a significant challenge for accurate pose reconstruction. This paper presents a benchmark on the robustness of 3D HPE models under realistic occlusion conditions, involving combinations of occluded keypoints commonly observed in real-world scenarios. We evaluate nine state-of-the-art 2D-to-3D HPE models, spanning convolutional, transformer-based, graph-based, and diffusion-based architectures, using the BlendMimic3D dataset, a synthetic dataset with ground-truth 2D/3D annotations and occlusion labels. All models were originally trained on Human3.6M and tested here without retraining to assess their generalization. We introduce a protocol that simulates occlusion by adding noise into 2D keypoints based on real detector behavior, and conduct both global and per-joint sensitivity analyses. Our findings reveal that all models exhibit notable performance degradation under occlusion, with diffusion-based models underperforming despite their stochastic nature. Additionally, a per-joint occlusion analysis identifies consistent vulnerability in distal joints (e.g., wrists, feet) across models. Overall, this work highlights critical limitations of current 3D HPE models in handling occlusions, and provides insights for improving real-world robustness.
Keywords 3D Human Pose Estimation  $\cdot$ Occlusions Robustness  $\cdot$ 2D Detections  $\cdot$ 2D-to-3D Lifting
1 Introduction
Human Pose Estimation (HPE) rom RGB images is a longstanding problem in computer vision field and plays a critical role in numerous real-world applications, including sports[1, 2, 3], healthcare[4, 5, 6, 7], and retail[8]. Specifically, 3D HPE involves estimating the three-dimensional coordinates of key human body points — such as elbows, knees, and wrists — from single images or video sequences. This capability allows systems to accurately interpret and analyze human movement, which is essential for tasks like motion capture in animation[9, 10, 11] and posture assessment in healthcare[5] or sports [3].
Figure 1: BlendMimic3D frames show a subject without (top left) and with occlusion (bottom left). Corresponding 3D ground-truth (top right) and simulated noisy prediction (bottom right) illustrate occlusion impact.
A central challenge in HPE is handling occlusions, where parts of the body are hidden from view—not only due to the subject’s own pose (self-occlusion), but also because of external objects or other people. For instance, in sports such as soccer or basketball, players often block one another, making it difficult for algorithms to reliably estimate keypoint locations[3]. In retail settings, limbs may be occluded by shelves or merchandise, complicating full-body pose detection. Similarly, in clinical gait analysis, assistive devices or medical equipment can obstruct key body parts[4, 6], requiring pose estimation methods to be robust under partial visibility.
Over the years, various models [12, 13, 14, 15, 16, 17, 18, 19, 20] have been developed to address this task, with a significant focus on the method of lifting from 2D to 3D poses. Despite recent advances, model performance continues to decline in the presence of occlusions, highlighting the critical need for improved robustness in 3D HPE systems.
Given this, our focus shifts to practical applications: How effectively can these models perform in real-world scenarios? To this end, we evaluate several state-of-the-art (SOTA) models for which code and pre-trained weights are publicly available [12, 13, 14, 15, 16, 17, 18, 19, 20].
In practical applications, two-stage models first capture a person on camera, detect 2D keypoints, and then lift them to a 3D pose. However, occlusions can lead to inaccuracies in detected 2D poses. As shown in Figure 1, these occlusion scenarios can lead to a decline in the performance of 3D reconstruction. This work focuses on 2D-to-3D lifting models and studies their ability to handle such inaccuracies: Can they effectively predict 3D poses from noisy 2D inputs?
Most existing studies use the Human3.6M [21] dataset, which is a controlled environment with indoor settings and minimal obstructions, except for occasional objects in scenarios such as sitting poses. This setup does not fully represent the challenges of real-world occlusions. We opt for the BlendMimic3D [22] dataset, which includes occlusion labels and synthetic images with occlusions, allowing us to specifically assess model performance under varying occlusion conditions. A challenge we faced was the disparity in keypoint formats across datasets [21, 22, 23]. Thus, we had to adjust the BlendMimic3D format to align it as closely as possible with Human3.6M, thereby facilitating the testing of SOTA models. We then introduce a benchmarking protocol that simulates varying occlusion severities by adding noise into the 2D keypoints labeled as occluded. This allows for systematic comparison of model performance under controlled occlusion conditions. Additionally, we perform per-keypoint sensitivity analyses to identify vulnerabilities in specific body regions.
Our findings reveal several key insights:
  • All models show significant degradation under occlusion, especially on distal keypoints (e.g., feet, wrists).
  • Diffusion-based models, despite their stochastic nature, underperform under occlusion unless equipped with explicit occlusion-aware designs.
  • Certain models (e.g., PoseFormerV2, DTF) show greater resilience due to architectural features such as frequency encoding or dual-view fusion.
  • There exists a critical gap between current benchmarks and real-world challenges, emphasizing the need for more occlusion-aware training and evaluation strategies in 3D HPE.
By analyzing model behavior under realistic occlusion and distribution shifts, this study aims to guide future research toward more robust pose estimation systems.
This paper is structured as follows. Section 2 reviews related work on 3D HPE, with a focus on occlusion handling. Section 3 describes the BlendMimic3D dataset, the selected models, the occlusion simulation protocols, and the evaluation metrics. Section 4 presents benchmarking results under varying occlusion conditions, including a per-keypoint occlusion analysis. Section 5 extends the analysis to the effects of dataset distribution shifts and camera placement. Section 6 discusses the broader challenges and limitations identified in current 3D HPE models. Finally, Section 7 concludes the paper and outlines directions for future research.
2 Related Work
Deep Learning has significantly advanced 2D Human Pose Estimation (HPE) [24, 25]. However, the same level of progress has not been achieved for 3D HPE, particularly in monocular settings, due to depth ambiguities, insufficient training data, and occlusions [22, 25, 26].
Recent surveys on 3D HPE have reviewed deep learning architectural trends [24], benchmark performances [27], and progress in monocular and multi-view setups [24, 28, 27]. However, these works typically evaluate performance using global accuracy metrics on controlled benchmarks like Human3.6M [21] or MPI-INF-3DHP [29], which do not reflect the complexities of real-world scenarios. Udayan et. al [26] recently conducted a systematic review of monocular 3D HPE, highlighting the gap between academic performance and practical applicability, and emphasizing the need to tackle challenges such as occlusions and depth ambiguity. However, their work is limited to a comparative synthesis of existing literature, they did not conduct new experiments to validate their conclusions.
Sarafinos et al.  [28] tackled the benchmark limitations by introducing a synthetic dataset that allows full control over scene variables and human pose configurations. Their goal was to test model robustness to anthropometric variability, viewpoint changes and action complexity. While this work offered valuable insights into classical methods, it primarily covered pre-deep learning techniques from 2008 to 2015 and included only a limited discussion of modern deep learning-based approaches. Building on this line of inquiry, our work extends the evaluation of occlusion robustness to current state-of-the-art (SOTA) deep learning models, leveraging a dataset specifically designed for fine-grained occlusion analysis. We begin by reviewing the main approaches to monocular 3D HPE, typically divided into direct estimation and 2D-to-3D lifting.
2.1 Direct 3D HPE
Direct methods infer 3D poses directly from RGB images without relying on intermediate 2D keypoint predictions. Li and Chan [30] proposed a dual-branch network leveraging body part correlations, demonstrating the potential of direct regression models. However, their approach lacked the precision and generalization required for complex poses or occluded scenarios.
Pavlakos et al. [31] introduced a voxel-based formulation that predicts 3D heatmaps for each keypoint, addressing depth ambiguity through volumetric likelihoods. However, their method requires post-processing to extract keypoint locations, which introduces non-differentiability and limits spatial precision. Moreover, achieving high resolution in voxel space increases memory and computation costs.
Sárándi et al. [32] tackled the issue of model generalization, arguing that deep learning-based 3D HPE improves when trained under a large amount of data. To address the challenge of merging training data from different datasets, which use inconsistent skeleton formats (i.e., label different anatomical landmarks), they introduce a latent keypoint set as an intermediate representation to unify the data for model training. While this improves training robustness, it does not explicitly enhance occlusion handling.
In contrast, Wang et. al [33] focused specifically on occlusion robustness by proposing Di2Pose, a diffusion-based model that predicts 3D poses as discrete tokens. Rather than relying on large-scale data unification, Di2Pose simulates occlude-and-replace transitions during training to implicitly model occlusions, enabling it to recover missing keypoint information without requiring explicit occlusion labels. Although their model enhance prediction accuracy under occlusions, its inference speed is limited by the iterative nature of diffusion sampling.
Despite these advancements, direct methods still face challenges in resolving depth ambiguities and occlusions due to the absence of intermediate spatial cues and limited contextual reasoning.
2.2 2D-to-3D HPE
Many studies have adopted 2D-to-3D lifting approaches, benefiting from the advances in 2D HPE. This two-stage approach has demonstrated superior performance on several benchmark datasets and is highly efficient, especially in settings where temporal context is available. Temporal information was incorporated by Pavllo et al. [12], they proposed the VideoPose3D model to address video sequences with a temporal dilated convolutional network. VideoPose3D is noted for its simplicity and efficiency. However, it assumes independence between consecutive errors, making it less robust to continuous occlusions or persistent noise.
Given that a human pose can be represented as a graph, where keypoints are nodes and bones are edges, graph-based approaches [34, 35, 36, 22] have been applied to model relationships between keypoints over time. Qiu et al. [34] introduced a Dynamic Graph Convolutional Network to capture temporal dependencies between 2D keypoints. Hu et al. [35] represented the 3D human skeleton as a directed graph to encode anatomical hierarchies, modeling bones as directed edges from parent to child keypoints, following the natural articulation from the hip (root) to the extremities. Cai et al. [36] proposed leveraging spatial-temporal relationships within a graph based framework. Building on these ideas, our previous work [22] introduced a 3D pose refinement block designed to correct occluded poses using graph-based reasoning.
While graph-based models have effectively exploited the structured nature of the human skeleton, capturing local and hierarchical dependencies among keypoints, transformer-based models have recently gained prominence due to their ability to capture long-range dependencies of keypoints. Transformer architectures [37, 13, 38, 39] process spatial-temporal features with global attention mechanisms, enabling keypoint reasoning across multiple pose sequences and leading to more robust 3D HPE. Zheng et al. [37] presented the first purely transformer-based 3D HPE approach, named PoseFormer, without convolutional architectures involved. This model processes both spatial and temporal aspects of human poses. Inspired by the progress of Masked Image Modeling (MIM) [40] in image classification, Shan et al. [13] proposed P-STMO, that applies masked keypoint modeling to 3D HPE with self-supervised learning. P-STMO improves accuracy in 3D pose estimation while reducing the need for large labeled datasets by randomly masking keypoints and frames using learnable embeddings, enabling the encoder to infer missing information. Zhang et al. [14] pointed out that PoseFormer overlooks motion differences between body keypoints and proposed MixSTE, a model that captures inter-frame correspondences using a mixed spatio-temporal encoder to extract fine-grained, keypoint-specific features. Building on PoseFormer, Zhao et al. [39] proposed PoseFormerV2, which leverages a compact frequency-domain representation of long skeleton sequences to efficiently scale up the receptive field and enhance robustness to noisy 2D keypoint detections, making it well-suited for complex and occluded scenarios.
Recently, diffusion-based models have emerged as a promising direction for 3D HPE [16, 14, 17, 18, 41], generating and refining multiple 3D hypotheses from 2D observations through a gradual denoising process. There models show potential in addressing challenges such as occlusions and depth ambiguity. Shan et al. [16] introduced D3DP, which generates multiple 3D pose candidates from a single 2D input, using a denoiser conditioned on 2D keypoints and adapted from MixSTE [14]. Choi et al. [17] proposed DiffuPose, which integrates a graph convolutional architecture to capture keypoint topology and guide the denoising process across the graph structure. Xu et al. [18] presented FinePose, a prompt-driven diffusion model that incorporates semantic information, such as body parts, action classes, and motion dynamics, to enable controllable pose generation, which shows promising results in handling complex poses and potential occluded scenarios.
Despite improved accuracy, diffusion models face challenges in computational efficiency due to their iterative nature, requiring multiple forward passes for each prediction. Moreover, their performance remains highly dependent on the quality of 2D inputs. Errors in 2D detection, especially under occlusions, can significantly degrade the final 3D estimation. In this work, we benchmark state-of-the-art models under varying levels of occlusion to assess their robustness and practical viability.
Even in the absence of occlusions, not all keypoints are equally difficult to estimate. In 2D detection, Yu et al. [42] showed that some keypoints are more easily detectable than others due to their scale. However, 3D HPE models typically allocate equal attention to all keypoints. In this work, we show that as the degrees of freedom increase, so does the complexity of 3D estimation for each keypoint, even in non-occluded scenarios.
2.3 Occlusion Handling in 3D HPE
Our previous work [22] employed graph-based 3D pose refinement block trained on occluded scenarios. While effective in these conditions, this approach does not consistently perform well in non-occluded settings, highlighting the need for occlusion-aware networks. Addressing this, Cheng et al. [43] included a cylinder man model to generate self-occlusion labels. Ghafoor and Mahmood [19] proposed T3D-CNN, an occlusion-guided framework based on temporal dilated convolutions, explicitly modeling missing keypoints via an occlusion guidance matrix. Later, Ghafoor et al. [20] extended this idea with a temporal interpolation-based occlusion guidance mechanism and a transformer-based approach that generates intermediate views, subsequently fused for final 3D pose estimation, a model named DTF. However, both T3D-CNN and DTF [19, 20] rely on the assumption that occlusion labels are known in advance. Their occlusion-guided frameworks depend on an external model to indicate which keypoints are occluded, treating these as missing inputs during prediction. Wang et al. [44] separate features into occlusion-aware and view-dependent components, completing occluded poses in 2D space, potentially leading to inaccurate 3D coordinates. Hardy and Kim [45] reconstruct occluded parts in 3D but do not address all occlusion types. Thus, there is still room for improvement in occlusion-aware models.
Despite recent advancements in the field [25], robust handling of occlusions under complex, real-world conditions remains a key open problem. To address this, our study evaluates the performance of several SOTA models under controlled occlusion scenarios using the BlendMimic3D dataset [22]. In this work, we leverage its existing occlusion labels to simulate 2D occlusion noise for comprehensive benchmarking.
2.4 3D HPE Generalization
Learning-based methods have achieved significant progress in 3D HPE, but their performance is highly dependent on the training datasets [32, 46]. In practice, models trained on one dataset often struggle when evaluated on others due to domain shifts and discrepancies in skeleton annotations. Jiang et al. [46] approached this challenge by integrating optimization strategies into a diffusion-based model that iteratively refines 3D pose predictions. However, because skeleton annotations differ across datasets, their diffusion model must be retrained for each dataset, limiting its generalization.
A common strategy to improve generalization is to train on multiple datasets. However, this introduces challenges related to differences in skeleton definitions [47, 32]. While many datasets use similarly named keypoints, their spatial definitions often differ—for instance, hip keypoints may be located at different heights, or some datasets label surface markers while others use anatomical interior points. Furthermore, certain keypoints may be missing entirely in some datasets. Mahmood et al. [48] introduced the AMASS dataset and addressed the problem of discrepancy in MoCap data representations by mapping them to the SMPL [49] representation, a parametric body mesh widely used in graphics and vision. However, since our focus lies in 3D skeleton-based pose prediction this solution is not directly applicable.
In 2D-to-3D pose lifting, Rapczyński et al. [47] proposed combining multiple datasets and manually aligning skeletons through hand-crafted rules (e.g., scaling the hip-to-pelvis distance). However, this approach scales poorly with large keypoint sets or many datasets. As an alternative, Sárándi et al. [32] proposed projecting poses from different datasets into a shared latent space via autoencoders. While effective as an intermediate format, this approach does not generate a definitive or universal keypoint schema.
2.5 2D and 3D HPE Datasets
Popular 2D HPE datasets include MPII [50], LSP [51], LSP-extended [52] and COCO  [23]. Among these, COCO has become the most widely used benchmark for image-based pose estimation due to its scale and diversity. However, being human-annotated, COCO is also prone to label noise and inconsistencies.
On the other hand, 3D HPE datasets typically require sophisticated equipment like motion capture (MoCap) systems for accurate pose recording [21, 29, 48, 53, 54, 55]. The most commonly used 3D HPE datasets include Human3.6M [21], one of the largest MoCap dataset with RGB video, 3D keypoint locations, and camera parameters. It is standard for benchmarking 3D pose estimation under controlled conditions. MPI-INF-3DHP [29] was captured with a commercial markerless MoCap system in both indoor and outdoor scenes, with multi-view RGB, green screen and studio backgrounds. HumanEva-I [54] is an early MoCap indoor dataset for monocular/multi-view 3D HPE. CMU Panoptic [55] is a multi-camera indoor dataset capturing social interactions between multiple people. While these datasets provide accurate 3D ground-truth, they often lack diversity, realism, and occlusion scenarios. To address these issues, both synthetic and in-the-wild datasets have emerged.
3DPW [56] is the first in-the-wild dataset with 3D ground-truth obtained via wearable IMUs and cameras in real-world environments, including multi-person settings. SURREAL [53] is a synthetic dataset with random combinations of 3D poses, body shapes, textures, and backgrounds. BEDLAM [57], a recent synthetic dataset designed for 3D human pose and shape (HPS) estimation, showed that models trained purely on synthetic data can generalize to real images. Other than these, recent efforts are moving toward more diverse (e.g., occlusion-heavy or multi-person[58]) and sensor-fused datasets.
Despite these recent advances, Human3.6M [21] remains the most widely used benchmark for single-person 3D pose, while COCO [23] dominates in the 2D HPE space. However, each has limitations, COCO’s human curated nature is prone to errors, and Human3.6M lacks occluded and real-world conditions.
Sárándi et al.[59] showed that SOTA 3D HPE models suffer significant performance drops when evaluated under synthetic occlusions. To address this, in our previous work we introduced BlendMimic3D [22], a synthetic dataset designed to include advanced occlusion scenarios. It includes action-specific sequences with keypoint visibility labels per frame, allowing precise evaluation under occlusion.
Most reviews focus on architectural efficiency or dataset diversity but lack a systematic evaluation of model performance under explicit occlusion conditions. In our study, we aim to bridge this gap by analyzing how SOTA 3D HPE models behave under different types and severities of occlusion, using BlendMimic3D dataset [22]. To the best of our knowledge, this is the first benchmark to perform a structured, occlusion-aware evaluation across a diverse set of models.
3 Experimental Setup
3.1 Selected Models
For this benchmark, we selected a diverse set of SOTA 2D-to-3D human pose estimation models covering different architectural types: convolutional, graph-based, transformer-based, and diffusion-based. Selection was based on the availability of open-source implementations, as well as their relevance and performance in the literature. All models were pretrained using an off-the-shelf 2D keypoint detector, CPN [60], on the Human3.6M [21] dataset, with a receptive field of 243 frames, except for DTF [20] and DiffuPose [17], which used 351 and 384 frames, respectively, as in their original proposals. Where pretrained weights were not available, we trained the models from scratch using authors’ official configurations, specifically for DiffuPose [17], DTF [20] and T3D-CNN [19]. All the selected models are summarized in Table 1. This selection enables a comprehensive comparison across methodologies, particularly under occlusion conditions explored in this study.
Table 1: Overview of the selected 2D-to-3D human pose estimation models benchmarked in this study. For each model, we summarize its main architectural focus, the availability of pretrained weights, and code references. Occlusion Aware Training (OAT).
Model Highlight OAT Pretrained Model
VideoPose3D [12, 61] Temporal convolutional network known for its simplicity and competitive performance.
P-STMO [13, 62] Transformer-based model trained with masked keypoint modeling for self-supervised learning.
MixSTE [14, 63] Mixed spatio-temporal transformer that captures fine-grained inter-keypoint motion dynamics.
PoseFormerV2 [15, 64] Frequency-domain transformer designed for long-sequence efficiency and robustness.
D3DP [16, 65] Diffusion-based model using a denoiser adapted from MixSTE.
DiffuPose [17, 66] Diffusion model with a graph convolutional backbone for topology-aware denoising.
FinePose [18, 67] Prompt-driven diffusion model incorporating action class, body part, and kinematic context.
T3D-CNN [19, 68] Occlusion-aware model using temporal convolutions guided by a keypoint-level occlusion mask.
DTF [20, 69] Occlusion-aware transformer-based model with view-fusion and occlusion interpolation.
For completeness, we note that our previously proposed model, a graph convolutional network (GCN) designed to refine 3D poses predicted from occluded 2D inputs [22], was not included in this evaluation. Since the model was trained directly on the BlendMimic3D dataset, using it in the same evaluation would compromise the validity of the benchmark.
Before analyzing model robustness under occlusions, we first evaluate the baseline performance of all selected models on the Human3.6M test set using 2D keypoints predicted by CPN [60]. This evaluation establishes an upper-bound reference for each model’s accuracy in an in-distribution setting. Table 2 summarizes the results.
Table 2: Baseline MPJPE of each model on Human3.6M using CPN detections. OAT denotes Occlusion-Aware Training. The best results are highlighted in bold and the second-best results are underlined.
Models Year OAT MPJPE - Avg [mm]
VideoPose3D [12] 2019 46.8
P-STMO [13] 2022 42.8
MixSTE [14] 2022 40.9
PoseFormerV2 [15] 2023 45.2
D3DP [16] (P-Agg) 2023 39.9
DiffuPose [17] (P-Agg) 2023 53.8
FinePose [18] (P-Agg) 2024 40.6
T3D-CNN [19](Rand 16) 2022 71.1
DTF-GT [20] (Rand 16) 2024 55.0
This value is higher than in their paper because they use HR-Net [70] as 2D pose detector. We retrained it with CPN detections.
This value is higher than in their paper because they use 2D ground-truth input. We retrained it with CPN detections.
For diffusion-based models, we report results using the Pose-level Aggregation (P-Agg) strategy, where multiple 3D pose hypotheses are averaged to produce a final output. We chose this approach because alternative aggregation strategies, such as JPMA [16], which relies on reprojection accuracy, can become unreliable in the presence of noisy or occluded 2D inputs. Additionally, methods like P-Best and J-Best, which require access to ground-truth 3D poses, are not feasible in real-world applications, where ground-truth 3D poses are unavailable at inference time.
For occlusion-aware models, namely T3D-CNN [19] and DTF [20], we replicate their training protocols using the Rand 16 setting, where 16 keypoints are randomly masked as occluded. These settings correspond to the configurations proposed by the authors in their original papers.
As shown in Table 2, D3DP achieves the best overall performance, followed closely by FinePose and MixSTE. These results highlight the effectiveness of diffusion-based models. Among transformer-based models, MixSTE, P-STMO and PoseFormerV2 all perform competitively, while VideoPose3D, despite its simplicity and earlier design, still offers a solid baseline. However, these numbers represent upper-bound performance, and do not reflect robustness to occlusion, addressed in the following sections.
On the other hand, occlusion-aware models T3D-CNN and DTF show higher MPJPE. This can be attributed to their training protocol, which simulates occlusion by randomly masking 16 keypoints in each frame (Rand 16), allowing the model to observe only one keypoint at a time. In this work, we evaluate how well these models generalize to real occlusion scenarios and whether this training scheme translates into improved robustness in practice.
3.2 Benchmark Dataset
To evaluate how each model generalizes to in-the-wild scenarios, we benchmark them on the BlendMimic3D [22] dataset. Since all selected models were trained on the Human3.6M [21] dataset, evaluating them on that same dataset would not assess their robustness to distribution shifts or unseen occlusions. BlendMimic3D provides a more realistic test set, with occlusions caused by object interactions and different viewpoints.
BlendMimic3D offers synchronized RGB videos from four camera views, along with ground-truth 2D and 3D pose annotations and occlusion labels, per keypoint and per frame. These occlusion annotations are crucial for this benchmark, as they allow us to simulate and manipulate different levels and types of realistic occlusion noise in a controlled manner.
We use subject S2 from BlendMimic3D, who performs 14 actions involving various occludes placed in the scene. To ensure compatibility across models and avoid retraining, we manually adapted the BlendMimic3D skeleton format in Blender to closely match the keypoint structure and conventions of Human3.6M, as can be seen in Figure 2. The converted 2D and 3D pose annotations, along with occlusion labels, will be made publicly available.
Figure 2: Comparison of 2D human pose representations across different formats. (Left) Human3.6M (H36M) format over its original image. (Middle) BlendMimic3D (BM3D) format and its adapted-to-H36M format and (Right) its respective poses.
In the Human3.6M dataset, keypoints such as Nose and Head are captured based on the orientation of the head using surrounding markers. Consequently, the position of the Nose or Head is not anatomically grounded and may vary across subjects and actions. In contrast, our Blender environment enforces anatomical consistency by fixing all keypoints, including Nose and Head, to precise, predefined locations.
Figure 3 shows the percentage of occlusion per keypoint across all frames, with an average of $33.93\%$ of keypoints occluded per frame. Notably, RHip and LHip are the most frequently occluded keypoints. However, in the Human3.6M format Hip, RHip and LHip are collinear (as illustrated in Figure 2). This geometric relationship allwos one to infer the third point if two are known. Our results confirm this: despite being consistently occluded, the hip keypoints are still well predicted by 3D HPE models. Figure 3 presents the distribution of occluded keypoints per frame. Notably, there are no frames where all keypoints are visible or completely occluded, the majority of frames have between 3 and 7 occluded keypoints.
Figure 3: Occlusion statistics of subject S2 in the BlendMimic3D dataset. (a) Percentage of occlusion per keypoint across all actions and frames. (b) Distribution of the number of occluded keypoints per frame.
3.3 Occlusion Simulation Protocol
In 2D-to-3D human pose estimation, the accuracy of the predicted 3D pose depends heavily on the quality of the 2D keypoints used as input. When body parts are occluded in the image, 2D keypoint detectors such as CPN [60] and Detectron2 [71], often produce noisy or inaccurate predictions. This makes occlusion in the 2D input space one of the primary sources of error for lifting-based models, and a critical challenge in real-world predictions.
To evaluate the robustness of 3D HPE models under occlusions, we designed two noise injection experiments using the BlendMimic3D dataset. In both cases, we simulate the effect of occlusion by corrupting 2D input poses with Gaussian noise, following the behavior of real 2D detectors under occluded conditions.
3.3.1 Protocol 1: Adding Noise to Occluded Keypoints
In order to understand how well each model can infer 3D poses from corrupted 2D input data, and thus assess the models’ practical usability in different occlusion levels, we tracked their performances across different noise levels. We applied zero-mean Gaussian noise with varying levels of standard deviation to occluded 2D keypoints, using the per-keypoint occlusion annotations available in BlendMimic3D. This approach allows us to systematically simulate uncertainty in 2D detections without introducing bias or directional shifts.
Let $\mathbf{x}_{ij}\in\mathbb{R}^{2}$ denote the 2D position of keypoint $j$ in frame $i$, and let $\mathcal{O}_{ij}\in\{0,1\}$ be the corresponding occlusion indicator, where $\mathcal{O}_{ij}=1$ denotes that the keypoint is occluded. We define the noise level parameter $\sigma\in\{0,0.001,0.005,0.01,0.03,0.05\}$, which is expressed as a fraction of the average image resolution. The corresponding standard deviation in pixel units is given by
$\sigma_{\text{px}}=\sigma\cdot\frac{w+h}{2},$
where $w$ and $h$ are the image width and height, respectively.
We inject noise into each occluded keypoint by sampling a 2D perturbation vector from an isotropic Gaussian distribution:
$\tilde{\mathbf{x}}_{ij}=\begin{cases}\mathbf{x}_{ij}+\boldsymbol{\varepsilon}_{ij},&\text{if }\mathcal{O}_{ij}=1\\ \mathbf{x}_{ij},&\text{otherwise}\end{cases}\quad\text{with}\quad\boldsymbol{\varepsilon}_{ij}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{px}}^{2}\mathbf{I}_{2})$
This process perturbs only the occluded keypoints in the 2D pose sequence, injecting Gaussian noise independently to the $x$ and $y$ coordinates of each occluded keypoint. We tested six noise levels, $\sigma=\{0,0.001,0.005,0.01,0.03,0.05\}$, applied over 10 independent runs per level for statistical robustness. The $\sigma=0$ case corresponds to ground-truth input. Figure 4 illustrates qualitative examples across noise levels. In our experiments we show that all models are highly sensitive to increased variance, even with zero mean, with notable degradation in 3D pose estimation accuracy.
Figure 4: 2D human poses derived from the ground-truth skeleton, with varying levels of Gaussian noise applied to the occluded keypoints: the nose, hands, knees, and feet.
3.3.2 Protocol 2: Per-Keypoint Occlusion Analysis
In the second evaluation, we analyze the sensitivity of each model to the occlusion of individual keypoints. The goal is to quantify the relative importance of each keypoint in accurate 3D pose reconstruction. For this, we selected the first 243 frames of each sequence and, over a 100-frame window (frame 50–149), for each of the 17 keypoints, injected zero-mean Gaussian noise into that keypoint alone, keeping all the others unaltered. This was repeated over five independent runs per keypoint to account for variability. By isolating occlusion to a single keypoint at a time, this test allows us to assess which keypoints impact 3D pose accuracy the most. The results help identify which body parts are easier or more difficult for each model to predict.
The standard deviation used for the Gaussian noise was chosen based on findings from the first evaluation and informed by real detector statistics, which are discussed later in the results section. The specific noise level is $\sigma=0.03$ and was introduced after analyzing the baseline sensitivity to ensure that it realistically reflects the magnitude of occlusion-induced errors observed in practice.
3.4 Evaluation Metrics
Our 3D human pose estimation evaluation is based on the Mean Per-keypoint Positional Error (MPJPE), which measures the average Euclidean distance (in millimeters) between the predicted and ground-truth 3D join positions. MPJPE is represented by the equation
$\text{MPJPE}=\frac{1}{N}\sum^{N}_{i=1}\lVert J_{i}-J_{i}^{*}\rVert_{2},$ (1)
where $N$ is the number of keypoints, and $J_{i}$, $J^{*}_{i}$, are the ground-truth and predicted positions of the $i_{th}$ keypoint, respectively.
To analyze model performance under occlusions, we compute and report the following variants:
  • Overall MPJPE: The average error over all keypoints in all frames, regardless of visibility.
  • Visible MPJPE: The average error computed over all keypoints labeled as visible across all frames. This isolates the model’s performance on clearly detected keypoints.
  • Occluded MPJPE: The average error computed over all keypoints labeled as occluded across all frames. This directly reflects the model’s ability to estimate uncertain keypoints.
All errors are scaled to millimeters and averaged over all actions and test runs for statistical robustness. When applicable, we report standard deviation as a confidence measure across runs.
Hip keypoint evaluation: We observed discrepancies in how different models handle the hip keypoint. Specifically, P-STMO, D3DP, FinePose and DTF, explicitly reset the predicted hip coordinates to zero before evaluation, regardless of their actual prediction. This post-processing step artificially reduces the MPJPE for the keypoint to zero, introducing an evaluation bias. To ensure consistency and fairness across models, we excluded the hip keypoint from all poses during evaluation.
4 Results and Discussion
4.1 Occlusion Severity
To assess how 3D HPE models handle increasing occlusion errors, we evaluated their performance using Protocol 1 (Section 3.3.1), where zero-mean Gaussian noise is applied to keypoints labeled as occluded in the BlendMimic3D annotations. We tested six levels of noise, $\sigma$, across ten independent runs. Table 3 reports MPJPE results (in millimeters) for each model.
Table 3: Evaluation of 3D HPE performance for state-of-the-art models under varying levels of simulated occlusion noise on the BlendMimic3D dataset. Results are reported as MPJPE (mm). The best results for each noise level are shown in bold, and the second-best are underlined. DTF-GT corresponds to the original configuration proposed in [20], where occluded keypoints are treated as missing (i.e., zero 2D coordinates) prior to interpolation. No missing keypoints, with simulated noise for the occluded keypoints. No interpolation, considering the ground-truth occlusion labels.
Noise Level
Models Year $\sigma=0$ $\sigma=0.001$ $\sigma=0.005$ $\sigma=0.01$ $\sigma=0.03$ $\sigma=0.05$
VideoPose3D [12] 2019 $73.94\pm 4.85$ $73.93\pm 4.69$ $74.09\pm 4.70$ $74.77$±$\pm 4.74$ $91.11$±$\pm 4.90$ $123.12\pm 7.99$
P-STMO [13] 2022 $75.44\pm 4.67$ $75.41\pm 4.52$ $74.95\pm 4.67$ $75.02\pm 4.93$ $114.58\pm 7.36$ $160.41\pm 9.87$
MixSTE [14] 2022 $\mathbf{69.46}\pm 4.61$ $\mathbf{69.51}\pm 4.50$ $71.76\pm 5.05$ $79.27\pm 6.68$ $131.94\pm 14.20$ $175.21\pm 17.03$
PoseFormerV2 [15] 2023 $73.20\pm 4.23$ $73.20\pm 4.10$ $73.28\pm 4.24$ $\mathbf{73.79}\pm 4.67$ $\mathbf{84.57}\pm 5.98$ $111.85\pm 8.28$
D3DP [16] 2023 $69.84\pm 3.74$ $69.81\pm 3.61$ $\mathbf{70.57}\pm 4.24$ $77.48\pm 5.97$ $127.82\pm 16.80$ $169.29\pm 23.58$
DiffuPose [17] 2023 $83.45\pm 6.99$ $83.66\pm 6.72$ $88.48\pm 6.69$ $101.20\pm 6.99$ $168.71\pm 12.68$ $223.32\pm 16.59$
FinePose [18] 2024 $69.75$±$\pm 8.87$ $69.78$±$\pm 8.65$ $70.78$±$\pm 9.67$ $77.34\pm 11.12$ $133.05\pm 18.84$ $183.86\pm 23.95$
DTF-GT [20] 2024 $80.83\pm 3.69$
DTF [20] 2024 $80.83\pm 3.69$ $93.37\pm 3.92$ $93.37\pm 3.95$ $93.44\pm 3.97$ $94.11\pm 4.24$ $\mathbf{95.70}\pm 5.02$
DTF [20] 2024 $80.83\pm 3.69$ $93.39\pm 3.92$ $93.37\pm 3.96$ $93.42\pm 4.01$ $94.14\pm 4.29$ $95.79$±$\pm 4.89$
Despite the noise applied to occluded keypoints being zero-mean, Table 3 reveals that all models exhibit clear sensitivity to increasing standard deviation. This indicates that the evaluated models are vulnerable to noise provided by 2D detectors when a keypoint is occluded.
The performance trends observed in Table 3 can be largely explained by the architectural designs of each model. VideoPose3D, a convolution-based baseline, maintains stable accuracy at low-to-moderate noise levels but lacks explicit mechanisms to handle severe occlusions, leading to a sharper decline under high noise.
P-STMO, despite its transformer-based design, shows limited robustness, likely due to its reliance on masked modeling strategies that may not generalize to real occlusion patterns. MixSTE, which alternates spatial and temporal transformer blocks, performs best in clean settings but degrades quickly under noise, suggesting its structured modeling may be sensitive to keypoint corruption. PoseFormerV2, in contrast, demonstrates strong resilience, especially at moderate noise levels, likely due to its use of frequency-domain features that help mitigate temporal inconsistencies.
D3DP, a diffusion-based method, maintains strong performance under moderate noise, highlighting the advantage of probabilistic modeling, though its accuracy decreases at higher noise levels. DiffuPose, while also diffusion-based, consistently underperforms, potentially due to fragile inference dynamics and limited noise conditioning. FinePose leverages fine-grained prompt conditioning, incorporating action and kinematic-aware guidance, and delivers competitive performance. However, it still lacks explicit mechanisms for occlusion robustness.
In conclusion, models like, MixSTE, D3DP, and PoseFormerV2 perform well under light to moderate occlusion noise, while DTF stands out under severe conditions. Yet none manage to maintain consistent performance across the full occlusion spectrum. This underscores a critical gap in the field: current state-of-the-art models struggle to balance accuracy on clean inputs with robustness to occlusion. Addressing this trade-off should be a central goal for future 3D human pose estimation research.
Although diffusion-based models are often expected to be robust to noisy inputs, this assumption does not fully hold here. All three tested methods, D3DP, DiffuPose, and FinePose, condition their denoising process directly on the 2D input pose, which guides every step. As a result, noise in occluded keypoints significantly degrades their predictions, revealing a consistent weakness in current diffusion designs when faced with realistic occlusions. Moreover, these models are computationally expensive, requiring iterative sampling during inference, which further limits their practicality in real-time.
[…]
As a complementary analysis to the results presented in Table 3, Figure 5 shows the normalized MPJPE, computed as the difference between each noise level and the baseline at $\sigma=0$, for three error types: overall pose error (average MPJPE), occluded keypoint error, and visible keypoint error. This visualization highlights how model performance degrades under increasing occlusion noise and reveals differences in sensitivity to visible versus occluded keypoints.
Figure 5: Normalized MPJPE across increasing noise levels. Each curve represents the MPJPE for overall, occluded, and visible keypoints, normalized by the respective model’s baseline performance at $\sigma=0$. (a) VideoPose3D, (b) P-STMO, (c) MixSTE, (d) PoseFormerv2, (e) D3DP, (f) DiffuPose, (g) FinePose, and (h) DTF.
Across all models, errors on occluded keypoints increase more steeply than on visible ones, revealing a common lack of explicit occlusion modeling. In contrast, DTF exhibits almost flat curves, with both visible and occluded MPJPE remaining stable across noise levels. MixSTE and FinePose show the greatest sensitivity to noise, particularly in occluded keypoints, where errors rise sharply related to visible keypoints.
DiffuPose suffers the most degradation across all error types, indicating poor robustness in its inference dynamics. D3DP maintains relative stability for visible keypoints but is notably affected by noise in occluded keypoints.
VideoPose3D, P-STMO and PoseFromerV2 remain relatively stable under mild noise ($\sigma<0.01$). Among them, P-STMO degrades faster, while VideoPose3D shows a more abrupt error increase as noise intensifies. PoseFormerV2 demonstrates the highest resilience, likely due to its frequency-domain temporal encoding, which helps mitigate the impact of occluded keypoint noise and supports more balanced performance.
4.1.1 DTF Performance
To accurately assess DTF’s behavior under realistic occlusion scenarios, we introduced three distinct test configurations detailed in Table 4. These correspond directly to the DTF results reported in Table 3. The original DTF model (DTF-GT) expects exactly 16 keypoints to be randomly occluded per frame. In their code, these occluded keypoints are set to zero (i.e., treated as missing), and their positions are then estimated through interpolation, guided by visibility-based confidence scores derived from temporal context. However, their assumption of having only one visible keypoint at a time and random occlusion patterns is completely unrealistic. In our evaluation using BlendMimic3D, which provides naturally occurring occlusion annotations, keypoints can remain occluded for longer durations than DTF expects, as can be seen in Figure 6. Such prolonged occlusions render DTF’s interpolation strategy less effective, indicating that its training approach is inadequate for real-world occlusions.
Table 4: DTF model configurations used in the occlusion robustness study, indicating use of occlusion labels, zeroed 2D keypoints, interpolation, and lifting. Variants help assess the role of each component.
Occl. Labels Zero 2D Coord. Interp. Lifting
DTF-GT
DTF
DTF
Figure 6: Analysis of the duration of occlusions per keypoint over the actions, from the ground-truth occlusion labels of BlendMimic3D subject S2, (a) the average duration and (b) maximum duration.
To bridge this gap between simulated and realistic occlusion, we evaluated two modified variants of the original DTF model: (1) DTF, which uses full 2D poses (including noisy occluded keypoints) and applies interpolation using the ground-truth occlusion labels, and (2) DTF†‡, which removes interpolation and directly inputs the noisy poses with occlusion labels. Table 4 summarizes the differences among these variants in terms of input preparation and processing.
[…]
On the other hand, configurations using raw noisy poses with occlusion labels (DTF, DTF†‡) yield more stable predictions, even as noise increases. Despite being designed for missing keypoints, DTF demonstrates robustness to real noisy poses, performing best at $\sigma=0.05$. This resilience stems from its dual-view generation and fusion architecture, along with a confidence-aware interpolation mechanism that enables the model to down-weight unreliable keypoints and leverage temporal consistency effectively.
We adopt the DTF configuration for all subsequent evaluations. This variant preserves the core components of the original model while using complete 2D poses, including occluded keypoints, rather than treating them as missing.
4.1.2 T3D-CNN Performance
Unlike DTF, T3D-CNN was designed to work with 2D keypoints masked as missing (set to zero), relying heavily on an internal occlusion mask provided as input. To evaluate its robustness, we trained T3D-CNN under three different masking strategies: no masking (None) , 8 randomly occluded keypoints (Rand 8) and 16 randomly occluded keypoints (Rand 16). Our goal was to assess the model’s ability to generalize across varying numbers of missing keypoints.
Table 5 presents the performance of each model under multiple test conditions: 2D ground-truth input from Human3.6M with no missing keypoints, with 8 or 16 randomly masked keypoints, and ground-truth 2D input from BlendMimic3D dataset, which includes realistic occlusions but lacks the original occlusion mask format expected by T3D-CNN.
[…]
Test Configuration
H36M
No miss Rand 8 Rand 16 BM3D
Model None 358.14
Rand 8 153.50 56.79 419.15 314.65
Rand 16 74.42
The results indicate that T3D-CNN struggles to generalize beyond the specific masking condition it was trained on. Performance degrades when tested on unseen masking configurations or on BM3D, where the number and duration of occluded keypoints vary naturally based on the scene and motion. Notably, the model trained with 8 randomly occluded keypoints (Rand 8) performs better on BM3D than the versions trained with no masking or 16 occlusions. This suggests that the configuration of Rand 8 is more reasonable and better aligned with the range of occlusion patterns found in BM3D, as shown in Figure 3. Because T3D-CNN cannot process complete or noisy 2D inputs, and lacks a mechanism to exploit real occlusion annotations, it was excluded from the noise-level evaluation presented in Table 3.
4.1.3 Occlusion Noise Analysis
After evaluating model robustness across a range of Gaussian noise levels (Protocol 1), we wanted to relate these simulations in real-world conditions. Specifically, we aimed to understand whether the selected noise levels realistically reflect the types and magnitudes of errors produced by standard 2D detectors under occlusion. Therefore, we conducted an error analysis of two commonly used 2D detectors, CPN [60] and Detectron2 [71], on 1000x1002 images from subject S2 in BlendMimic3D. The detectors output 2D poses in COCO format [23], which we aligned semantically with the Human3.6M-converted ground-truth, excluding non-aligned keypoints such as head, neck, spine, eyes and ears. Table 6 presents the MPJPE and standard deviation for both visible and occluded keypoints. Occlusions were further categorized into two groups: self-occlusions, where a subject occludes themselves, and external occlusions, caused by objects or out-of-frame positioning.
Table 6: Mean and standard deviation of 2D pose errors (MPJPE in pixels) for visible and occluded keypoints (Kpts) using CPN and Detectron2 on subject S2 of BlendMimic3D.
CPN [60] Detectron2 [71]
Mean Std Mean Std
Visible Kpts 6.11 3.53 6.84 4.21
Self-Occluded Kpts 11.47 4.72 14.27 6.28
Externally-Occluded Kpts 19.77 16.65 28.07 23.74
[…]
Additionally, the results in Table 3 and Figure 5 reinforce this categorization. For $\sigma=\{0.001,0.005\}$, many models maintain or slightly improve their performance compared to using ground-truth 2D input. This suggests that such low levels of noise reflect the standard deviation typically present in training with detected 2D poses. At $\sigma=0.01$, the noise magnitude begins to simulate mild occlusion behavior while still partially overlapping with the visible range. Once the noise increases to $\sigma=0.03$, the performance of the models drops more noticeably, reflecting a realistic degree of occlusion-induced perturbation.
From this analysis, we conclude that zero-mean Gaussian noise with $\sigma=0.03$ provides a reliable simulation of real occlusions in 2D keypoints, and will be used in the subsequent per-keypoint occlusion evaluation (Protocol 2).
4.2 Per-Keypoint Occlusion Analysis
In the second part of our study, we focus on identifying which keypoints are more vulnerable to occlusion and whether this behavior persists when the same keypoints are visible. Building on the detector noise and occlusion severity analyses (Sections 4.1 and 4.1.3), we selected $\sigma=0.03$ as a representative occlusion noise level, simulating moderate to severe occlusions.
To isolate the impact of each keypoint, in each experiment, we added zero-mean Gaussian noise into one individual 2D keypoint, between frames 50 and 149 of the first 243-frame segment. This localized occlusion simulation allows us to evaluate per-keypoint error propagation in the 3D predictions. We refer to this procedure as Protocol 2, described in Section 3.3.2.
Figure 7 displays the aggregated MPJPE results across all models. We distinguish the error on the occluded keypoint, the remaining visible keypoints, and the full-pose average. From Figure 7 we can see that a common pattern emerges: certain keypoints are inherently more difficult to estimate when occluded. Notably, the wrists (Lwrist, Rwrist), feet (RFoot, LFoot), Nose and Head show higher MPJPE under occlusion. These extremities are small, fast-moving, and often heavily occluded, making them particularly error-prone. In contrast, more central and structurally constrained keypoints, such as the hips, knees, shoulders, and elbows, exhibit lower errors, even under occlusion.
Figure 7: MPJPE per keypoint across occluded frames, where each keypoint is occluded independently using Gaussian noise ($\sigma=0.03$). We report (i) MPJPE of the occluded keypoint, (ii) MPJPE of visible keypoints, and (iii) average pose error. Values are averaged across all models.
To understand whether this error pattern is purely due to occlusion or an intrinsic modeling difficulty, we compare occluded keypoint errors with their errors under fully visible conditions. Figure 8 shows that even when all keypoints are visible, these same keypoints (e.g., wrists, feet) still exhibit higher MPJPE, indicating an inherent estimation difficulty.
Figure 8: Comparison of MPJPE per keypoint when all keypoints are visible (2D ground truth) versus when each keypoint is individually occluded. Values are averaged across all models.
We extended this analysis to real occlusions in the BlendMimic3D dataset using the provided labels and added Gaussian noise ($\sigma=0.03$). In Figure 9, we show the MPJPE of each keypoint when visible vs. occluded within real occlusion scenarios. The same pattern persists: error increases with occlusion, and the most affected keypoints remain the same. This supports our conclusion that keypoints with higher degrees of freedom (knees, feet, wrists, and elbows) are consistently harder to estimate.
[ht] Refer to caption
Figure 9: MPJPE for each keypoint under real occlusion conditions in BlendMimic3D (BM3D). We compare errors when the keypoint is visible versus occluded across poses. These reflect the natural distribution of occlusions and keypoint visibility. Values are averaged across all models.
A detailed per-model breakdown of all three analyses, per-keypoint occlusion (Figure 7), comparison with fully visible pose (Figure 8), and real occlusion analysis (Figure 9), is provided in the supplementary material. These results include model-specific plots of occluded vs. visible keypoint errors, enabling a deeper understanding of how each architecture responds to occlusion.
5 Beyond Occlusions: Evaluating Additional Factors
While our primary analysis focused on model robustness under varying levels of occlusions, it is important to consider other factors that might also contribute to performance degradation. In this section, we explore how dataset distribution shifts, particularly related to camera-subject configuration, can affect model generalization. All evaluated models were trained on Human3.6M [21], a highly controlled indoor dataset, whereas our benchmark uses BlendMimic3D [22], a synthetic dataset designed to simulate more realistic and diverse occlusion scenarios. We also briefly mention other possible factors, such as velocity and action diversity, which were found to have little impact (see supplementary material).
5.1 Domain Shift: Human3.6M vs. BlendMimic3D
[…]
Figure 10: Data distribution of subject orientation and distance to the camera in Human3.6M (a) and BlendMimic3D (b).
5.2 Camera-Related Sensitivity
To investigate whether model errors correlate with subject orientation and distance to the camera, we analyzed performance heatmaps on BlendMimic3D using ground-truth 2D poses (i.e., without occlusions or detector noise). Figure 11 shows the average MPJPE across all models, aggregated over bins of orientation and distance.
Figure 11: Average MPJPE across all models as a function of subject orientation and distance to the camera. Evaluated on BlendMimic3D with ground-truth 2D inputs.
We observe a consistent trend: performance degrades at shorter distances (below  2.5 meters), a range underrepresented in Human3.6M (see Figure 10). This highlights a generalization challenge due to distribution shift in camera-subject configurations. Beyond this, orientation and distance do not exhibit a clear correlation with MPJPE when results are averaged across models. For a more detailed understanding of model-specific sensitivity patterns, including how individual architectures react to different orientations and distances, per-model heatmaps can be found in supplementary material.
We also analyzed velocity and action variability, and concluded that they do not influence our results. This analysis is provided in the supplementary material.
6 Challenges and Limitations
Throughout this study, several challenges and limitations of current 3D HPE models have emerged.
Skeleton format mismatch is a key issue. Most 3D HPE models are trained using the Human3.6M keypoint format, whereas real-world 2D detectors often output COCO-format keypoints. This mismatch can introduce errors unless a careful alignment and mapping are performed between the two formats. For our benchmark, we manually aligned the BlendMimic3D dkeleton format to match Human3.6M using Blender software.
We also identified an evaluation bias related to hip keypoint centering. Many models normalize poses by subtracting the 3D coordinates of the hip, effectively zeroing the hip and removing the global trajectory during training. However, some models, including P-STMO, D3DP, FinePose, and DTF, explicitly reset the predicted hip to zero during evaluation, regardless of the actual output. This post-processing step can artificially reduce MPJPE and introduces inconsistency across models. To ensure fairness, we excluded the hip keypoint from evaluation for all models.
One of the most critical limitations is occlusion handling. Most state-of-the-art models suffer a noticeable drop in performance when keypoints are partially or fully occluded. In particular, models like D3DP, despite showing strong results under clean conditions, tend to degrade rapidly as occlusion levels increase. This indicates that robustness to occluded inputs remains an open challenge in 3D HPE.
We also found that model performance varies significantly across body parts. Keypoints with greater degrees of freedom, such as the hands, feet, and head, consistently resulted in higher errors, especially under occlusion. This points to intrinsic difficulties in modeling highly dynamic keypoints.
Camera-related distribution shifts pose another challenge. Our analysis showed that models trained solely on Human3.6M struggle when evaluated under camera distances or viewing angles that were not present during training, particularly at closer distances. This reveals a clear gap in the generalization ability of current approaches.
Taken together, these findings show that despite strong performance on benchmark datasets, current 3D HPE models are not yet robust for real-world applications, especially in scenarios involving occlusion, diverse camera setups, and dynamic poses.
7 Conclusion
In this study, we conducted a comprehensive benchmark of state-of-the-art 2D-to-3D human pose estimation models under realistic occlusion scenarios using the BlendMimic3D dataset. Our focus was to evaluate model robustness to noisy 2D inputs, a common challenge in real-world applications.
To this end, we introduced two controlled evaluation protocols. The first simulated increasing occlusion severity by adding Gaussian noise into all occluded keypoints based on ground-truth occlusion labels. The second evaluated the impact of occluding each keypoint individually, allowing us to isolate per-keypoint vulnerabilities.
Our evaluation revealed several key findings. All models, regardless of architecture, suffer performance degradation under occlusion. Diffusion-based models, despite their probabilistic formulation, proved sensitive to noisy inputs since their denoising processes rely heavily on the 2D input pose. Occlusion-aware models such as DTF showed improved robustness under severe occlusion but underperformed on clean inputs. Protocol 2 revealed that keypoints with higher degrees of freedom, such as wrists and feet, are consistently more error-prone, even when not occluded.
Beyond occlusions, we highlighted that camera-related distribution shifts between training (Human3.6M) and evaluation (BlendMimic3D) affect generalization, especially at closer distances and unseen viewpoints. We also identified evaluation inconsistencies, such as hip keypoint normalization and skeleton format mismatches, which we addressed to ensure fair comparison across models.
Despite advances in 3D HPE, our results show that current models are not yet robust enough for real-world environments. Future research should focus on developing occlusion-aware models that not only generalize across diverse viewing conditions and skeleton formats, but also maintain strong performance under clean input settings.
Appendix A Per-Keypoint Occlusion Analysis (Per Model)
This appendix complements the main per-keypoint occlusion analysis by providing detailed, model-specific results. While Figures 7, 8, and 9 in the main paper show average performance across all models, here we present the corresponding individual results for each method under three complementary conditions:
  • Synthetic occlusion of one keypoint at a time, with analysis of the MPJPE for the occluded keypoint, the remaining visible keypoints, and the full-pose average (Figure 12).
  • Comparison of per-keypoint MPJPE when all keypoints are visible versus when one is occluded (Figure 13).
  • Realistic occlusion scenarios using ground-truth occlusion labels from BlendMimic3D (Figure 14).
Across all figures, MPJPE is computed for each keypoint and reported separately for: (i) occluded keypoints, (ii) visible keypoints, and (iii) average pose error. The shaded regions represent standard deviations across sequences.
Figure 12: MPJPE per keypoint across occluded frames, where each keypoint is occluded independently using Gaussian noise ($\sigma=0.03$). We report (i) MPJPE of the occluded keypoint, (ii) MPJPE of visible keypoints, and (iii) average pose error. Models: (a) VideoPose3D, (b) P-STMO, (c) MixSTE, (d) PoseFormerv2, (e) D3DP, (f) DiffuPose, (g) FinePose, and (h) DTF.
From Figure 12, we observe that MixSTE and FinePose are the most sensitive models to occlusions, followed by P-STMO, DiffuPose and D3DP. This is reflected in the higher MPJPE of occluded keypoints compared to visible ones. A similar behavior is evident in Figure 13, where the error difference between occluded and visible settings is more pronounced for the same group of models. In contrast, models such as VideoPose3D, PoseFormerV2, DTF show minimal differences between occluded and visible errors per keypoint (Figure 13), suggesting greater robustness to localized perturbations.
Figure 13: Comparison of MPJPE per keypoint when all keypoints are visible (GT 2D) vs. when each keypoint is individually occluded with Gaussian noise. Models: (a) VideoPose3D, (b) P-STMO, (c) MixSTE, (d) PoseFormerv2, (e) D3DP, (f) DiffuPose, (g) FinePose, and (h) DTF.
Figure 14 shows that, under real occlusions most models reproduce the patterns seen in the per-keypoint case, confirming that keypoints like wrists, feet, and head remain the most error-prone. D3DP and DiffuPose, although diffusion-based, degrade significantly under real occlusions, revealing vulnerabilities in their conditioning mechanisms.
Figure 14: MPJPE for each keypoint under real occlusion conditions in BlendMimic3D (BM3D). We compare errors when the keypoint is visible versus occluded across poses. These reflect the natural distribution of occlusions and keypoint visibility. Models: (a) VideoPose3D, (b) P-STMO, (c) MixSTE, (d) PoseFormerv2, (e) D3DP, (f) DiffuPose, (g) FinePose, and (h) DTF.
Together, these results support the aggregated trends discussed in the main paper and reinforce the importance of occlusion-aware training strategies.
Appendix B Beyond Occlusions: Camera-Related Sensitivity (Per Model)
While our primary analysis focused on robustness to occlusions, it is important to consider other contextual factors that may influence model performance, such as camera-related aspects. This appendix complements the camera-related sensitivity analysis presented in the main paper by providing detailed, model-specific heatmaps. While Figure 11 in the main paper reports average MPJPE across all models, here we examine individual model performance as a function of subject orientation and distance to the camera (see Figure 15).
Figure 15: Model performance as a function of subject orientation and distance from the camera, evaluated on BlendMimic3D under ground-truth 2D inputs. Models: (a) VideoPose3D, (b) P-STMO, (c) MixSTE, (d) PoseFormerv2, (e) D3DP, (f) DiffuPose, (g) FinePose, and (h) DTF.
Across all models, we observe a consistent trend: performance degrades at closer distances (below approximately 2.5 meters). As discussed in the main paper, this is likely due to the lack of close-range examples in the training data (see Figure 10).
Beyond this consistent degradation at shorter distances, no clear or systematic trend emerges with respect to subject orientation or longer-range distance. Each model exhibits distinct and often inconsistent performance patterns across orientation angles. While some models, such as VideoPose3D, DiffuPose, and DTF, maintain relatively stable errors across orientations, others, including P-STMO, MixSTE, PoseFormerV2, D3DP, and FinePose display scattered performance spikes concentrated in different regions. This variability suggests orientation and distance do not affect model performance in a predictable way.
Appendix C Beyond Occlusions: Action Variability (Per Model)
In addition to camera-related analysis, we also investigate the impact of action variability. Different actions include varying poses and spatial configurations, which may affect prediction accuracy. Figure 16 reports the MPJPE for each action across all camera viewpoints, averaged across all models, while Figure 17 presents the corresponding per model performances. These figures show that the test set includes a broad and realistic distribution of actions and viewpoints (see also Figure 10), supporting the diversity needed to evaluate generalization.
Figure 16: Average MPJPE across all models for each action and camera, under no occlusions.
Figure 17: Per-model MPJPE for each action under no occlusions. Results are shown for four camera viewpoints and the average across views. Models: (a) VideoPose3D, (b) P-STMO, (c) MixSTE, (d) PoseFormerv2, (e) D3DP, (f) DiffuPose, (g) FinePose, and (h) DTF.
From Figure 16 we can conclude that there is no single action that consistently degrades performance across all cameras or models. While the performance of Camera 0 shows slightly more variability across models, overall performance remains relatively stable across actions in the absence of occlusions.
To further assess action-related variability, we evaluate the per-model curves in Figure 17. No consistent performance trend is observed across actions. Each model displays a unique performance profile depending on the camera and action. However, one exception is the SittingCheering action, which frequently appears among the worst-performing across models, likely due to its complex, asymmetric movements. These findings suggest that action variability alone does not strongly influence model performance in the same way occlusions do.
Appendix D Beyond Occlusions: Pose Velocity Analysis (Per Model)
Pose velocity, how fast keypoints move between frames, can also challenge temporal models, especially if velocity distributions in the test set differ from those seen during training. Figure 18 compares the distribution of 2D pose velocities (in pixels per frame) between Human3.6M (training set) and BlendMimic3D (test set), both on a logarithmic scale. While the two datasets differ slightly in distribution, the overlap suggests that test-time velocities are within the training domain.
Figure 18: Velocity distribution (in log scale) of keypoint motion in Human3.6M (top) and BlendMimic3D (bottom).
Figure 19 provides a per-model breakdown of model performance as a function of keypoint velocity. Each plot shows the frequency of keypoint observations across velocity bins, along with the corresponding MPJPE. All models except DiffuPose show a small increase in error with velocity. Interestingly, DiffuPose shows higher errors at lower velocities, improving as velocity increases. This counterintuitive behavior may indicate that the diffusion-based model tends to over-process clean inputs, attempting to denoise even when unnecessary.
Figure 19: MPJPE as a function of keypoint velocity for each model. Frequency of keypoint occurrences is shown in the background. Models: (a) VideoPose3D, (b) P-STMO, (c) MixSTE, (d) PoseFormerv2, (e) D3DP, (f) DiffuPose, (g) FinePose, and (h) DTF.
Overall, the slight initial error increase followed by near-stable performance across velocities suggests that velocity has minimal impact on MPJPE. These results confirm that velocity differences do not account for the performance degradation observed under occlusion, reinforcing the specificity of our occlusion sensitivity analysis.